Find z by the use of determinant:
3x-4y+2z+8=0
x+5y-3z+2=0
5x+3y-z+6=0
1
Expert's answer
2015-09-08T07:53:38-0400
Find z by the use of determinant:
3x-4y+2z+8=0
x+5y-3z+2=0
5x+3y-z+6=0

Solution

According Cramer's rule

| 3 -4 2|
D = det| 1 5 -3|= 120
| 5 3 -1|

| 3 -4 -8|
Dz =det | 1 5 -2| = 24
| 5 3 -6|


z=Dz/D = 5